OpenStudy (leenathan):

What is the volume of a sphere with a diameter of 14 centimeters? Use 3.14 for pi. Round your answer to the nearest hundredth. 1077.02 cm3 1436.03 cm3 1975.43 cm3 11,494.04 cm3

OpenStudy (leenathan):

@StudyGurl14

OpenStudy (leenathan):

@sissyedgar

OpenStudy (studygurl14):

Volum eof Sphere \(\Large\frac{4}{3}\pi r^3\)

OpenStudy (leenathan):

i think its b

OpenStudy (studygurl14):

Remember that r = 1/2 diamter

OpenStudy (studygurl14):

correct

OpenStudy (leenathan):

thanks

OpenStudy (leenathan):

i have more ill post more can you help with those?